Let's say that a company's profitability, p(t), follows p(t)= t^2 - 5t - 50 with
p(t) in units of $1K and t = fiscal months. How long in years will it take the company to be $100K in the black (i.e., p(t)=$100,000)?
a. 15 years
b. 1 1/2 years
c. 1 1/4 years
d. 1/2 year

Answer by ankor@dixie-net.com(22740)

: Let's say that a company's profitability, p(t), follows p(t)= t^2 - 5t - 50 with
p(t) in units of $1K and t = fiscal months. How long in years will it take the company to be $100K in the black (i.e., p(t)=$100,000)?
:
The equation for this problem, solve for t:
t^2 - 5t - 50 = 100
:
t^2 - 5t - 50 - 100 = 0
:
t^2 - 5t - 150 = 0
:
Factor this to:
(t-15)(t+10) = 0
:
The positive solution:
t = 15 yrs to be 100 k in the black